OpenStudy (anonymous):

Solve for x: 3x- (x-9) = 2(x+5)

jimthompson5910 (jim_thompson5910):

Start off by distributing

jimthompson5910 (jim_thompson5910):

It may help to think of 3x- (x-9) = 2(x+5) as 3x- 1(x-9) = 2(x+5)

OpenStudy (anonymous):

yeah i did but i got 2x-9=2x+10 so i don't know what the answer is because no number can be x. would it be null?

jimthompson5910 (jim_thompson5910):

2x-9=2x+10 2x-9-2x=2x+10-2x 0x-9= 0x+10 0-9 = 0+10 -9 = 10... which is FALSE So there are no solutions (due to the contradiction in the last equation)

OpenStudy (anonymous):

okay thanks!

jimthompson5910 (jim_thompson5910):

np